OpenStudy (anonymous):

Differentiate 8/(X+2)

OpenStudy (anonymous):

f(x) = 8/u(x) df(x)/dx=[(df/du)](du/dx) = [-8/u(x)^2)](1) I think that is how it's done. "Chain rule"?

OpenStudy (raffle_snaffle):

Let y=8/(x+2) so